Write two data sets with a range equal to 35 but with different standard deviations. You do not need to calculate the standard deviations.
Answer by jim_thompson5910(35256)

Data Set #1: 1, 5, 33, 36

Range = 36 - 1 = 35
Standard Deviation is relatively large (since the data values are spread out)



Data Set #2: 1, 2, 4, 7, 9, 12, 13, 14, 16, 19, 20, 21, 25, 27, 29, 30, 31, 34, 35, 36

Range = 36 - 1 = 35
Now the standard deviation is slightly smaller (since the data values are closer together)